1.HOW MANY OUNCES OF PEANUTS SHOULD BE ADDED TO 38OZ TO MAKE 4 POUNDS OF MIXED NUTS.
2.by compounding interest semiannually at the annual rate of 12 percent each $1 is worth $2.01219 at the end of 6 years. find the total interest earned of $1,000 compounded semiannually at 12 percent annual rate at the end of 6 years.
i have no idea with that one

Answer by stanbon(75887)

1. 8 pounds = 16(8)=128 ounces
128-32=90 ounces
Therefore 90 ounces must be add to the 32 to get 8 pounds.
2. A=P(1+r/n)^(nt)
A=1000(1+0.12/2)^(2(6))
A=1000(1.06)^12
A=1000(2.01219...*
A=$2012.19
Cheers,
Stan H.